OpenStudy (anonymous):

What is the equation of the line, in slope-intercept form, that passes through point (-2, 5) and is parallel to the line whose equation is x + y = 10? A.)y=x- 3 B.)y=x+ 3 C.)y= -x+ 3

OpenStudy (tyteen4a03):

Hint: The slope-intercept equation is formed like this: \(y = mx + c\), where m is the slope and c is the y-intercept. 2 lines that are parallel to each other have the same slope.

OpenStudy (anonymous):

Okay, the thing is, is that im awful at anything to do with algebra so I REALLY need you to explain it to me.

OpenStudy (tyteen4a03):

Whoops, you should use the point-slope form instead: y – y1 = m(x – x1), where y1 and x1 are the x and y coordinates of the coord pair, and m is the slope. Firstly, find the slope of the equation x + y = 10. Then, substitute the values in. Lastly, turn it into the slope-intercept form by algebra, and you have your answer.